use the fundamental theorem of integral calculus to evaluate the integral integration of 0 to 1 ( 2xsinx 1/x - cos.1/x )dx



1
Expert's answer
2021-11-02T11:50:06-0400
"\displaystyle\int_{0}^{1}(2x\sin(1/x)-\cos(1/x))dx"

Let "t=x^2\sin(1/x)." Then


"dt=(2x\sin(1/x)-\cos(1/x))dx"

"\int(2x\sin(1/x)-\cos(1/x))dx=\int dt=t+C"

"=x^2\sin(1/x)+C"

"\displaystyle\int_{0}^{1}(2x\sin(1/x)-\cos(1/x))dx"

"=\lim\limits_{A\to0^+}\displaystyle\int_{A}^{1}(2x\sin(1/x)-\cos(1/x))dx"


"=\lim\limits_{A\to0^+}[x^2\sin(1/x)]\begin{matrix}
 1 \\
 A
\end{matrix}"

"=1^2\sin(1/1)-\lim\limits_{A\to0^+}(A^2\sin(1/A))"

Use the Squeeze Theorem


"-A^2\leq A^2\sin(1/A)\leq A^2, A\in \R"

"\lim\limits_{A\to0^+}(-A^2)=\lim\limits_{A\to0^+}(A^2)=0"

Then by the Squeeze Theorem


"\lim\limits_{A\to0^+}(A^2\sin(1/A))=0"

Therefore


"\displaystyle\int_{0}^{1}(2x\sin(1/x)-\cos(1/x))dx=\sin(1)"